# Wrong Answers Can Also Be Useful: PlausibleQA — A Large-Scale QA Dataset with Answer Plausibility Scores

### v1

@@ Abstract.
Dataset, Question Answering, Large Language Models
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; July 13–18, 2025; Padova, ITisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Evaluation of retrieval resultsccs: Information systems Test collections
@@ 3.1.3. Question Sampling
Assume that you are unaware that the answer to $<$question$>$ is $<$ground_truth$>$. Generate a list of 10 unique candidate answers, ensuring that $<$ground_truth$>$ is excluded. A plausibility score evaluates how reasonable, credible, or contextually appropriate each candidate answer is in relation to the given question. For each candidate, provide: (1) A non-zero plausibility score as a number between 0 and 100. (2) A detailed explanation of the reasoning behind the plausibility score. Format your response as a JSON list, where each candidate is represented as:
@@ 3.2. Candidate Answer Generation
To generate candidate answers, we use LLaMA-3.3-70B (Grattafiori et al., 2024) as the LLM, as it is a highly capable model with strong performance in various NLP tasks. Also, it is open-source and available on HuggingFace [^4] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct, enabling users to reimplement our proposed pipeline and generate candidate answers for their custom datasets.
@@ 4.2. Human Evaluation
You are a knowledgeable assistant. Answer the following multiple-choice question by selecting the most appropriate option. Question: $<$question$>$ A) $<$option_1$>$ B) $<$option_2$>$ C) $<$option_3$>$ D) $<$option_4$>$ The letter of the chosen option must be from [A, B, C, D]. Format your response as a JSON item, where each output is represented as:
@@ 5.1. MCQA Results
The purpose of using difficult (Hard-wA and Hard-woA) and easy (Easy-wA and Easy-woA) options is to assess the accuracy of plausibility scores and their ability to challenge MCQA systems in identifying the correct answer from the options. Additionally, comparing conditions where the correct answer is present in the options (Hard-wA and Easy-woA) versus cases where None of them is included (Hard-woA and Easy-woA) helps evaluate how the presence of the correct answer impacts MCQA system performance.
@@ 6. Limitations and Use Cases
  • The dataset could support research in automatic hint generation and evaluation, particularly for assessing the Convergence metric (Mozafari et al., 2024b) which evaluates how well the hints narrow down potential answers to the question.

### v2

@@ Abstract.
Dataset, Question Answering, Multi-Choice QA, Robustness, Large Language Models
journalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730299isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Evaluation of retrieval resultsccs: Information systems Test collections
@@ 3.1.3. Question Sampling
Assume that you are unaware that the answer to $<$question$>$ is $<$ground_truth$>$. Generate a list of 10 unique candidate answers, ensuring that $<$ground_truth$>$ is excluded. A plausibility score evaluates how reasonable, credible, or contextually appropriate each candidate answer is in relation to the given question. For each candidate, provide: (1) A non-zero plausibility score as a number between 0 and 100. (2) A detailed explanation of the reasoning behind the plausibility score. Format your response as a JSON list, where each candidate is represented as: [
@@ 3.2. Candidate Answer Generation
To generate candidate answers and compare pairs, we use LLaMA-3.3-70B (Grattafiori et al., 2024) as the LLM, as it is a highly capable model with strong performance in various NLP tasks. Also, it is open-source and available on HuggingFace [^4] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct, enabling users to reimplement our proposed pipeline and generate candidate answers for their custom datasets.
@@ 4.2. Human Evaluation
[…]
You are a knowledgeable assistant. Answer the following multiple-choice question by selecting the most appropriate option. Question: <<<question>>> A) <<<option_1>>> B) <<<option_2>>> C) <<<option_3>>> D) <<<option_4>>> The letter of the chosen option must be from [A, B, C, D]. Format your response as a JSON item, where each output is represented as: {
@@ 5.1. MCQA Results
The purpose of using difficult (Hard-wA and Hard-woA) and easy (Easy-wA and Easy-woA) options is to assess the accuracy of plausibility scores and their ability to challenge MCQA systems in identifying the correct answer from the options. Additionally, comparing conditions where the correct answer is present in the options (Hard-wA and Easy-wA) versus cases where None of them is included (Hard-woA and Easy-woA) helps evaluate how the presence of the correct answer impacts MCQA system performance.
@@ 6. Limitations and Use Cases
  • The dataset could support research in automatic hint generation and evaluation, particularly for assessing the Convergence metric (Mozafari et al., 2025), which measures how effectively hints narrow down potential answers to a question. The resulting scores could contribute to a more accurate assessment of convergence within the HintEval framework [^11] https://github.com/DataScienceUIBK/HintEval.